Please write down Faraday’s Law and Ampere-Maxwell’s equation for time-dependent electromagnetic fields in a medium.


1
Expert's answer
2021-12-24T11:32:41-0500

Faraday’s Law:


"\nabla\times\textbf E=-\frac{\partial\textbf B}{\partial t}."

Ampere-Maxwell’s equation:


"\nabla\times\textbf H-(\sigma+i\omega\epsilon)\textbf E=0."